A bag contains 6 red, 5 green, and 7 white balls. The probability of choosing a red or a white ball is
A. 1/3 B. 11/13 C. 13/18 D. 3/8

Grade:Upto college level

1 Answers

Latika Leekha
9 years ago
The bag contains 6 red, 5 green and 7 white balls.
Probability of selecting a particular ball = Number of that balls/ (Total number of balls)
The probability of selecting a red ball i.e. P(Red) = 6/ 18
The probability of selecting a green ball i.e. P(Green) = 5/ 18
The probability of selecting a white ball i.e. P(White) = 7/ 18
Hence P (Red or White ball) = P(Red) + P(White)
= 6/18 + 7/18
= 13/18.
Thanks & Regards
Latika Leekha
askIITians Faculty